A line through (3, -4) with a slope of -9 has a y-intercept at ??
Answer by lwsshak3(11628)

A line through (3, -4) with a slope of -9 has a y-intercept at ??
***
Standard form of equation for a line: y=mx+b, m=slope, b=y-intercept
Equation: y=-9x+b
solve for b using coordinates of given point (3,-4)
-4=-9*3+b
-4=-27+b
b=-23
Equation: y=-9x+23
y-intercept:(0, 23)